The Aleutian Company produces two products, Rings and Dings. They are manufactured in two departments, Fabrication and assembly. Data for the products and departments are listed below.

                                 number of                 Labor hrs                    Machine hours

Product                             units                   per unit                           per unit

Rings                               1000                            4                                    6

Dings                                2000                            3                                   9


All of the machine hours take place in the Fabrication department, which has an estimated overhead of $90,000. All of the labor hours take place in the Assembly department, which has an estimated total overhead of $105,000.

The Aleutian Company uses a departmental overhead rates. The fabrication department uses machine hours for an allocation base, and the assembly department uses labor hours.


What is the overhead cost per unit for Dings?
a. $56.75

b. $65.25

c. $23.25

d. $64.50
Answer #1
Total Lab Hrs = LH pu*Rings + LHpu*Dings = 4*1000 +3*2000 = 10000 LHs
Total Assly Dept OH = $105,000
So Lab Hr OH rate = $105,000/10000 = $10.50

Total MHs = MH pu*Rings + MHpu*Dings = 6*1000 +9*2000 = 24000 MHs
Total Fab Dept OH = $90,000
So Mc Hr OH rate = $90,000/24000 = $3.75

So Overhead cost per unit for Dings = LH Rate*No of Lab Hrs + MH Rate* No of MHs
= $10.50*3 + $3.75*9 = $65.25 .....Ans (b)